### v2

# Second version: August 2024 Third version: November 2024 This version: September 2025)

Abstract
We revisit size controllability results in Pötscher and Preinerstorfer (2025) concerning heteroskedasticity robust test statistics in regression models. For the special, but important, case of testing a single restriction (e.g., a zero restriction on a single coefficient), we povide a necessary and sufficient condition for size controllability, whereas the condition in Pötscher and Preinerstorfer (2025) is, in general, only sufficient (even in the case of testing a single restriction).
1 Introduction
Tests and confidence intervals based on so-called heteroskedasticity robust standard errors date back to Eicker (1963, 1967) and constitute, at least since White (1980), a major component of the applied econometrician’s toolbox. Although these early methods come with well-understood large sample properties, when based on critical values derived from asymptotic theory their finite sample properties often deviate substantially from what asymptotic theory suggests: tests may substantially overreject under the null and corresponding confidence intervals may undercover. Strong leverage points have been identified early on as one major reason for these deviations, see, e.g., MacKinnon and White (1985), Davidson and MacKinnon (1985), and Chesher and Jewitt (1987). This has led to various developments trying to attenuate such drawbacks:
  1. 1.
    modifications of the covariance matrix estimators in Eicker (1963, 1967) and White (1980) led to tests based on what are now frequently called HC1-HC4 covariance matrix estimators (see, e.g., Long and Ervin (2000), and Cribari-Neto (2004) for an overview of the relevant literature), with HC0 denoting the original proposal;
  2. 2.
    some authors investigated degree-of-freedom corrections to obtain modified critical values (e.g., Satterthwaite (1946) or Bell and McCaffrey (2002), see also Imbens and Kolesár (2016));
  3. 3.
    wild bootstrap methods were investigated (for an overview of the relevant literature see Pötscher and Preinerstorfer (2023)) and, more recently, parametric bootstrap methods were studied in Chu et al. (2021) and Hansen (2021).
Although these developments sometimes lead to improvements, they come with no general finite sample guarantees concerning the size of the tests or the coverage of related confidence intervals, cf. the discussion in Pötscher and Preinerstorfer (2023, 2025) for detailed accounts.
Motivated by this lack of finite sample guarantees, Pötscher and Preinerstorfer (2025) studied the question under which conditions heteroskedasticity robust test statistics as well as the standard (uncorrected) F-test statistic can actually be paired with appropriate (finite) critical values, so that one obtains tests that have their (finite sample) size controlled by the prescribed significance value $\alpha$ (i.e., have size $\leq\alpha$) even though one is completely agnostic about the form of heteroskedasticity. [^1] The null-hypothesis to be tested is given by a set of affine restrictions. Under appropriate assumptions on the errors, allowing for Gaussian as well as substantial non-Gaussian behavior, they have shown that the standard (uncorrected) F-test statistic can be size-controlled (in finite samples) by using an appropriately chosen (finite) critical value if and only if the following simple condition holds:
no standard basis vector that lies in the column span of the design matrix
is “involved” in the affine restrictions to be tested, (1)
see (8) in Pötscher and Preinerstorfer (2025) for a formal statement of this condition.
Under a generally stronger condition than (1) (see (10) in Pötscher and Preinerstorfer (2025)), it was furthermore shown that large classes of heteroskedasticity robust test statistics (e.g., HC0-HC4) can be size-controlled by appropriate (finite) critical values. That condition, however, although satisfied for many testing problems (and even often identical to (1), cf. Theorem 3.9 and Lemma A.3 in Pötscher and Preinerstorfer (2018)), is not necessary in general, as shown in examples given in Pötscher and Preinerstorfer (2025); e.g., their Example 5.5 or Example C.1 in their Appendix C. [^2] Appendices to Pötscher and Preinerstorfer (2025) are published in the Supplementary Material available at the publisher’s website of that article. These examples consider the case of testing linear contrasts in the expected outcomes of subjects belonging to two or more groups, scenarios that are practically relevant. Further examples are provided in Examples A.1-A.4 in Appendix A further below. [^3] Example 5.5 in Pötscher and Preinerstorfer (2025) concerns simultaneously testing multiple retrictions, while Example C.1 in Appendix C of Pötscher and Preinerstorfer (2025) as well as Examples A.1-A.4 in Appendix A of the present article concern the case of testing a single restriction.
For the important case of testing problems involving only a single restriction (i.e., the case $q=1$ in the notation of Pötscher and Preinerstorfer (2025)), we show in the present article that the condition in (1) is then in fact necessary and sufficient also for size controllability of the above mentioned classes of heteroskedasticity robust test statistics, including HC0-HC4.
2 Results on size controllability
2.1 Framework
Here we recall the most relevant notions from Sections 2 and 3 of Pötscher and Preinerstorfer (2025), to which we refer the reader for further information and discussion. We consider the linear regression model
$\mathbf{Y}=X\beta+\mathbf{U},$ (2)
where $X$ is a (real) nonstochastic regressor (design) matrix of dimension $n\times k$ and where $\beta\in\mathbb{R}^{k}$ denotes the unknown regression parameter vector. Throughout, we assume $\mathop{\mathrm{r}ank}(X)=k$ and $1\leq k<n$. We furthermore assume that the $n\times 1$ disturbance vector $\mathbf{U}=(\mathbf{u}_{1},\ldots,\mathbf{u}_{n})^{\prime}$ ( denoting transposition) has mean zero and unknown covariance matrix $\sigma^{2}\Sigma$ ($0<\sigma<\infty$), where $\Sigma$ varies in the “heteroskedasticity model” given by
$\mathfrak{C}_{Het}=\left\{\mathop{\mathrm{d}iag}(\tau_{1}^{2},\ldots,\tau_{n}^{2}):\tau_{i}^{2}>0\text{ for all }i\text{, }\sum_{i=1}^{n}\tau_{i}^{2}=1\right\},$
and where $\mathop{\mathrm{d}iag}(\tau_{1}^{2},\ldots,\tau_{n}^{2})$ denotes the diagonal $n\times n$ matrix with diagonal elements given by $\tau_{i}^{2}$. That is, the disturbances are uncorrelated but can be heteroskedastic of arbitrary form. [In Appendix B we shall also consider another heteroskedasticity model.] [^4] Since we are concerned with finite-sample results only, the elements of $\mathbf{Y}$, $X$, and $\mathbf{U}$ (and even the probability space supporting $\mathbf{Y}$ and $\mathbf{U}$) may depend on sample size $n$, but this will not be expressed in the notation. Furthermore, the obvious dependence of $\mathfrak{C}_{Het}$ on $n$ will also not be shown in the notation, and the same applies to the heteroskedasticity model defined in Appendix B.
For ease of exposition, we shall maintain in the sequel that the disturbance vector $\mathbf{U}$ is normally distributed. Generalizations to classes of non-normal disturbances can be obtained following the arguments in Section 7.1 of Pötscher and Preinerstorfer (2025), see Remark 2.2 further below. Denoting a Gaussian probability measure with mean $\mu\in\mathbb{R}^{n}$ and (possibly singular) covariance matrix $A$ by $P_{\mu,A}$, the collection of distributions on $\mathbb{R}^{n}$ (the sample space of $\mathbf{Y}$) induced by the linear model just described together with the Gaussianity assumption is then given by
$\left\{P_{\mu,\sigma^{2}\Sigma}:\mu\in\mathrm{\mathop{\mathrm{s}pan}}(X),0<\sigma^{2}<\infty,\Sigma\in\mathfrak{C}_{Het}\right\},$
where $\mathrm{\mathop{\mathrm{s}pan}}(X)$ denotes the column space of XX. [^5] Since every $\Sigma\in\mathfrak{C}_{Het}$ is positive definite, the measure $P_{\mu,\sigma^{2}\Sigma}$ is absolutely continuous with respect to Lebesgue measure on $\mathbb{R}^{n}$.
We focus on testing the null $R\beta=r$ against the alternative $R\beta\neq r$, where $R\neq 0$ is a $1\times k$ vector and $r\in\mathbb{R}$. That is, throughout this paper we focus on testing a single restriction, whereas the theory developed in Pötscher and Preinerstorfer (2025) allows for simultaneously testing multiple restrictions (that is, we here consider only the special case corresponding to $q=1$ in Pötscher and Preinerstorfer (2025)). Set $\mathfrak{M}=\mathop{\mathrm{s}pan}(X)$, define the affine space
$\mathfrak{M}_{0}=\left\{\mu\in\mathfrak{M}:\mu=X\beta\text{ and }R\beta=r\right\},$
and let
$\mathfrak{M}_{1}=\left\{\mu\in\mathfrak{M}:\mu=X\beta\text{ and }R\beta\neq r\right\}.$
Adopting these definitions, the testing problem we consider can be written more precisely as
$H_{0}:\mu\in\mathfrak{M}_{0},\ 0<\sigma^{2}<\infty,\ \Sigma\in\mathfrak{C}_{Het}\quad\text{ vs. }\quad H_{1}:\mu\in\mathfrak{M}_{1},\ 0<\sigma^{2}<\infty,\ \Sigma\in\mathfrak{C}_{Het}.$ (3)
We also write $\mathfrak{M}_{0}^{lin}=\mathfrak{M}_{0}-\mu_{0}=\left\{X\beta:R\beta=0\right\}$ where $\mu_{0}\in\mathfrak{M}_{0}$. Of course, $\mathfrak{M}_{0}^{lin}$ does not depend on the choice of $\mu_{0}\in\mathfrak{M}_{0}$. Furthermore, if $\mathcal{L}$ is a linear subspace of $\mathbb{R}^{n}$, $\Pi_{\mathcal{L}}$ denotes the orthogonal projection onto $\mathcal{L}$, while $\mathcal{L}^{\bot}$ denotes the orthogonal complement of $\mathcal{L}$ in $\mathbb{R}^{n}$.
The assumption of nonstochastic regressors made above entails little loss of generality, and results for models with stochastic regressors can be obtained from the ones derived in the present paper by the same arguments as the ones given in Section 7.2 of Pötscher and Preinerstorfer (2025).
2.2 Test statistics, size controllability, and a new result
We consider the same test statistics as in Section 3 of Pötscher and Preinerstorfer (2025). Simplified to the setting of testing a single restriction considered in the present article, they are given by
$T_{Het}\left(y\right)=\left\{\begin{array}[]{cc}(R\hat{\beta}\left(y\right)-r)^{2}/\hat{\Omega}_{Het}(y)&\text{if }\hat{\Omega}_{Het}\left(y\right)\neq 0,\\ 0&\text{if }\hat{\Omega}_{Het}\left(y\right)=0,\end{array}\right$. (4)
where $\hat{\beta}(y)=\left(X^{\prime}X\right)^{-1}X^{\prime}y$ and where $\hat{\Omega}_{Het}(y)=R\hat{\Psi}_{Het}(y)R^{\prime}$. Here
$\hat{\Psi}_{Het}\left(y\right)=(X^{\prime}X)^{-1}X^{\prime}\mathop{\mathrm{d}iag}\left(d_{1}\hat{u}_{1}^{2}\left(y\right),\ldots,d_{n}\hat{u}_{n}^{2}\left(y\right)\right)X(X^{\prime}X)^{-1},$
with $\hat{u}(y)=\left(\hat{u}_{1}(y),\ldots,\hat{u}_{n}(y)\right)^{\prime}=y-X\hat{\beta}(y)$. The constants $d_{i}>0$ sometimes depend on the design matrix; see Pötscher and Preinerstorfer (2025) for examples of the weights $d_{i}$, including HC0-HC4 weights. We also recall the following assumption from the latter reference, again specialized to the setting of testing only a single restriction (i.e., to the case $q=1$ in the notation of Pötscher and Preinerstorfer (2025)).
Assumption 1.
Let $1\leq i_{1}<\ldots<i_{s}\leq n$ denote all the indices for which $e_{i_{j}}(n)\in\mathop{\mathrm{s}pan}(X)$ holds where $e_{j}(n)$ denotes the $j$-th standard basis vector in $\mathbb{R}^{n}$. If no such index exists, set $s=0$. Let $X^{\prime}\left(\lnot(i_{1},\ldots i_{s})\right)$ denote the matrix which is obtained from $X^{\prime}$ by deleting all columns with indices $i_{j}$, $1\leq i_{1}<\ldots<i_{s}\leq n$ (if $s=0$, no column is deleted). Then $R(X^{\prime}X)^{-1}X^{\prime}\left(\lnot(i_{1},\ldots i_{s})\right)\neq 0$ holds.
This assumption can be checked in any particular application as it only depends on the observable quantities $R$ and $X$; and a sufficient condition for Assumption 1 obviously is $s=0$. Assumption 1 is unavoidable if one wants to obtain a sensible test from the statistic $T_{Het}$, see Section 3 of Pötscher and Preinerstorfer (2025) for more discussion. We note that $e_{j}(n)\in\mathop{\mathrm{s}pan}(X)$ is equivalent to $h_{jj}=1$, where $h_{jj}$ denotes the $j$-th diagonal element of the ‘hat matrix’ $H=X(X^{\prime}X)^{-1}X^{\prime}$. [^6] This follows from $h_{jj}=e_{j}(n)^{\prime}He_{j}(n)=(He_{j}(n))^{\prime}He_{j}(n)$ and the fact that $H$ represents the orthogonal projection onto $\mathop{\mathrm{s}pan}(X)$.
As in Pötscher and Preinerstorfer (2025), we introduce
$B(y)=R(X^{\prime}X)^{-1}X^{\prime}\mathop{\mathrm{d}iag}\left(\hat{u}_{1}(y),\ldots,\hat{u}_{n}(y)\right).$
Define (recall that $R$ is a nonzero row vector in this article)
$\mathsf{B}=\left\{y\in\mathbb{R}^{n}:\mathop{\mathrm{r}ank}(B(y))<1\right\}=\left\{y\in\mathbb{R}^{n}:B(y)=0\right\}.$
It is now easy to see that $\mathop{\mathrm{s}pan}(X)\subseteq\mathsf{B}$ and that $\mathsf{B}$ is a linear space (cf. also Lemma 3.1 in Pötscher and Preinerstorfer (2025)). Simple examples can be constructed to show that $\mathop{\mathrm{s}pan}(X)\neq\mathsf{B}$, in general, cf. Example C.1 in Appendix C of Pötscher and Preinerstorfer (2025) as well as Examples A.1-A.4 in Appendix A further below.
To summarize the main size controllability statements from Pötscher and Preinerstorfer (2025) for the above class of test statistics, we first have to recall the following notation: For a given linear subspace $\mathcal{L}$ of $\mathbb{R}^{n}$ we define the set of indices $I_{0}(\mathcal{L})$ via
$I_{0}(\mathcal{L})=\left\{i:1\leq i\leq n,e_{i}(n)\in\mathcal{L}\right\}.$ (5)
We set $I_{1}(\mathcal{L})=\left\{1,\ldots,n\right\}\backslash I_{0}(\mathcal{L})$. Clearly, $\mathop{\mathrm{c}ard}(I_{0}(\mathcal{L}))\leq\dim(\mathcal{L})$ holds. And $I_{1}(\mathcal{L})$ is nonempty provided $\dim(\mathcal{L})<n$; in particular, $I_{1}(\mathfrak{M}_{0}^{lin})$ is always nonempty since $\dim(\mathfrak{M}_{0}^{lin})=k-1<n-1$. The results in Pötscher and Preinerstorfer (2025) concerning size controllability of tests for (3) based on $T_{Het}$ can now be summarized as follows; some intuition for why size control cannot always be achieved is provided further below as well as in Section 4 in Pötscher and Preinerstorfer (2025):
Theorem 2.1 (Theorem 5.1(b,c) and Propositions 5.5(b) and 5.7(b) in Pötscher and Preinerstorfer (2025) for the case $q=1$).
 [^7] The corresponding results in Pötscher and Preinerstorfer (2025) for $q\geq 1$ take exactly the same form, but with the definitions of the relevant quantities adapted to that more general setting.
Suppose that Assumption 1 is satisfied. Then the following statements hold:
  1. 1.
    For every $0<\alpha<1$ there exists a real number $C(\alpha)$ such that
    $\sup_{\mu_{0}\in\mathfrak{M}_{0}}\sup_{0<\sigma^{2}<\infty}\sup_{\Sigma\in\mathfrak{C}_{Het}}P_{\mu_{0},\sigma^{2}\Sigma}(T_{Het}\geq C(\alpha))\leq\alpha$ (6)
    holds, provided that
    $e_{i}(n)\notin\mathsf{B}\text{ \ \ for every \ }i\in I_{1}(\mathfrak{M}_{0}^{lin}).$ (7)
    Furthermore, under condition (7), even equality can be achieved in (6) by a proper choice of $C(\alpha)$, provided $\alpha\in(0,\alpha^{\ast}]\cap(0,1)$ holds, where
    $\alpha^{\ast}=\sup_{C\in(C^{\ast},\infty)}\sup_{\Sigma\in\mathfrak{C}_{Het}}P_{\mu_{0},\Sigma}(T_{Het}\geq C)$ (8)
    is positive and where
    $C^{\ast}=\max\{T_{Het}(\mu_{0}+e_{i}(n)):i\in I_{1}(\mathfrak{M}_{0}^{lin})\}$ (9)
    for $\mu_{0}\in\mathfrak{M}_{0}$ (with neither $\alpha^{\ast}$ nor $C^{\ast}$ depending on the choice of $\mu_{0}\in\mathfrak{M}_{0}$).
  2. 2.
    Suppose (7) is satisfied. Then a smallest critical value, denoted by $C_{\Diamond}(\alpha)$, satisfying (6) exists for every $0<\alpha<1$. And $C_{\Diamond}(\alpha)$ is also the smallest among the critical values leading to equality in (6) whenever such critical values exist.
  3. 3.
    Suppose (7) is satisfied. Then any $C(\alpha)$ satisfying (6) necessarily has to satisfy $C(\alpha)\geq C^{\ast}$. In fact, for any $C<C^{\ast}$ we have $\sup_{\Sigma\in\mathfrak{C}_{Het}}P_{\mu_{0},\sigma^{2}\Sigma}(T_{Het}\geq C)=1$ for every $\mu_{0}\in\mathfrak{M}_{0}$ and every $\sigma^{2}\in(0,\infty)$.
  4. 4.
    If the condition
    $e_{i}(n)\notin\mathop{\mathrm{s}pan}(X)\text{ \ \ for every \ }i\in I_{1}(\mathfrak{M}_{0}^{lin})$ (10)
    is violated, then $\sup_{\Sigma\in\mathfrak{C}_{Het}}P_{\mu_{0},\sigma^{2}\Sigma}(T_{Het}\geq C)=1$ for every choice of critical value $C$, every $\mu_{0}\in\mathfrak{M}_{0}$, and every $\sigma^{2}\in(0,\infty)$ (implying that size equals $1$ for every CC). [^8] It is understood here that critical values are less than infinity.
Note that condition (10) can equivalently be expressed in terms of certain diagonal elements of the ‘hat matrix’ $H$, see (12) in Remark 2.1 below. This is, however, not the case for condition (7). An informal verbal description of (10) is given in (1) in the Introduction.
To obtain some intuition for Theorem 2.1, recall that the diagonal elements of $\Sigma\in\mathfrak{C}_{Het}$ are positive and sum up to one (by definition). Now, for a matrix $\Sigma$ with $i$-th diagonal entry close to $1$, all other diagonal entries must therefore be close to 0, so that $\Sigma\approx e_{i}(n)e_{i}(n)^{\prime}$ then holds. Note that if $\Sigma\approx e_{i}(n)e_{i}(n)^{\prime}$, the distribution $P_{\mu,\sigma^{2}\Sigma}$ of the data is strongly “concentrated” around the one-dimensional space $\mu+\mathop{\mathrm{s}pan}(e_{i}(n))$. From an intuitive point of view, whether a given test statistic admits a size-controlling critical value or not, should therefore depend on the “behavior” of the test statistic for values on or close to the spaces $\mu_{0}+\mathop{\mathrm{s}pan}(e_{i}(n))$ with $\mu_{0}\in\mathfrak{M}_{0}$. It turns out that this is intimately related to (7) and (10). See Section 4 in Pötscher and Preinerstorfer (2025) for more discussion.
Most importantly, the above theorem shows that, given Assumption 1, the condition in (7) is sufficient for the existence of a (finite) size-controlling critical value $C(\alpha)$ satisfying (6), while the weaker condition (10) is necessary. Furthermore, in case the design matrix $X$ and the vector $R$ are such that $\mathsf{B}=\mathop{\mathrm{s}pan}(X)$, and hence the condition in (7) coincides with that in (10), the condition (7) is also necessary. However, $\mathsf{B}=\mathop{\mathrm{s}pan}(X)$ is not always true (see Example C.1 in Appendix C of Pötscher and Preinerstorfer (2025) or the examples in Appendix A further below), although the equality holds generically (cf. Theorem 3.9 and Lemma A.3 in Pötscher and Preinerstorfer (2018)). We now show in the subsequent theorem that in the situation considered in this article, namely testing only a single restriction, the condition in (7) in Theorem 2.1 can actually be replaced by that in (10).
Theorem 2.2.
Suppose that Assumption 1 is satisfied. Then the following statements hold:
  1. 1.
    For every $0<\alpha<1$ there exists a real number $C(\alpha)$ such that (6) holds, provided that (10) holds. Furthermore, under condition (10), even equality can be achieved in (6) by a proper choice of $C(\alpha)$, provided $\alpha\in(0,\alpha^{\ast}]\cap(0,1)$ holds, where $\alpha^{\ast}$ given by (8) is positive and where $C^{\ast}$is given by (9) for $\mu_{0}\in\mathfrak{M}_{0}$ (with neither $\alpha^{\ast}$ nor $C^{\ast}$ depending on the choice of $\mu_{0}\in\mathfrak{M}_{0}$).
  2. 2.
    Suppose (10) is satisfied. Then a smallest critical value, denoted by $C_{\Diamond}(\alpha)$, satisfying (6) exists for every $0<\alpha<1$. And $C_{\Diamond}(\alpha)$ is also the smallest among the critical values leading to equality in (6) whenever such critical values exist.
  3. 3.
    Suppose (10) is satisfied. Then any $C(\alpha)$ satisfying (6) necessarily has to satisfy $C(\alpha)\geq C^{\ast}$. In fact, for any $C<C^{\ast}$ we have $\sup_{\Sigma\in\mathfrak{C}_{Het}}P_{\mu_{0},\sigma^{2}\Sigma}(T_{Het}\geq C)=1$ for every $\mu_{0}\in\mathfrak{M}_{0}$ and every $\sigma^{2}\in(0,\infty)$.
  4. 4.
    If (10) is violated, then $\sup_{\Sigma\in\mathfrak{C}_{Het}}P_{\mu_{0},\sigma^{2}\Sigma}(T_{Het}\geq C)=1$ for every choice of critical value $C$, every $\mu_{0}\in\mathfrak{M}_{0}$, and every $\sigma^{2}\in(0,\infty)$ (implying that size equals $1$ for every CC). [^9] Cf. Footnote 8.
The main take-away of Theorem 2.2 is that given Assumption 1 holds, the condition in (10) is necessary and sufficient for the existence of a (smallest) finite size-controlling critical value when one is testing only a single restriction. Note that the conditions in (7) and (10) do not depend on the weights used in the construction of the covariance matrix estimator or on $r$. They only depend on $X$ and $R$. This and more (e.g., how the conditions relate to high-leverage points) is discussed subsequent to Theorem 5.1 (and in Remarks 5.2-5.4, 5.6, and 5.9) in Pötscher and Preinerstorfer (2025) to which we refer the reader for a detailed account. As a point of interest we also note that condition (10) given above is exactly the same as condition (8) in Pötscher and Preinerstorfer (2025) (with $q=1$); in that reference, the latter condition is shown to be necessary and sufficient for size control of the standard (uncorrected) F-test statistic (regardless of whether $q=1$ or not).
We also note here that Theorem 2.2 disproves – for the special case of testing a single restriction – a conjecture in Remark 5.8 of Pötscher and Preinerstorfer (2025), namely that there would exist cases where Assumption 1 holds, (10) is satisfied, (7) does not hold, and size control by a (finite) critical value is not possible.
To see why the refinement of Theorem 2.1 provided in Theorem 2.2 can matter in practice, it is enough to consider the textbook example of a matrix $X$ with two columns, the first indicating membership to the treatment group and the second indicating membership to the control group (a special case of Example C.1 in Appendix C of Pötscher and Preinerstorfer (2025)). Assume that the first $n_{1}\geq 2$ observations belong to the treatment group and the remaining $n_{2}\geq 2$ observations belong to the control group. Assume further that one wants to test whether $\beta_{1}$, the expected outcome of the treatment group, equals a given value (e.g., because one wants to obtain a confidence interval through test inversion). Example C.1 in Appendix C of Pötscher and Preinerstorfer (2025) shows that in this case
$\mathcal{I}_{1}(\mathfrak{M}_{0}^{lin})=\{1,\ldots,n\}\quad\text{ and }\quad\mathsf{B}=\{y\in\mathbb{R}^{n}:y_{1}=\ldots=y_{n_{1}}\}\neq\mathop{\mathrm{s}pan}(X).$
In particular, $e_{i}(n)\in\mathsf{B}$ if and only if $i>n_{1}$, so that (7) is not satisfied, while (10) holds, and size-controlling critical values hence exist by Theorem 2.2 (and can be used for constructing confidence intervals). Further examples are provided in Appendix A below.
Remark 2.1: (i) Condition (10) is equivalent to $h_{ii}<1$ for every $i\in I_{1}(\mathfrak{M}_{0}^{lin})$”. [^10] Note that $h_{ii}=1$ always holds if $i\in I_{0}(\mathfrak{M}_{0}^{lin})$.
(ii) Condition (10) can also equivalently be written as
$e_{i}(n)\notin\mathop{\mathrm{s}pan}(X)\text{ for every }i\text{ satisfying }R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}\neq 0,$ (11)
see Remark B.1(iii) in Appendix B further below. [^11] Comparing (10) and (11) could lead one to conjecture equivalence of the conditions $i\in I_{1}(\mathfrak{M}_{0}^{lin})$ and $R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}\neq 0$. This is incorrect in general, see Example A.1 in Appendix A. However, $R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}\neq 0$ implies $i\in I_{1}(\mathfrak{M}_{0}^{lin})$, see Part 3 of Lemma A.4. And this in turn is now equivalent to
$h_{ii}<1\text{ for every }i\text{ satisfying }R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}\neq 0.$ (12)
The last form of the condition may be more appealing to some readers. We issue a warning here, however, namely that the condition (7) is, in general, stronger than the condition $e_{i}(n)\notin\mathsf{B}$ for every $i$ satisfying $R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}\neq 0$”, see Remark B.1(iv) in Appendix B.
(iii) The condition $e_{i}(n)\notin\mathop{\mathrm{s}pan}(X)$ for every $i=1,\ldots,n$” (which is tantamount to $h_{ii}<1$ for every $i=1,\ldots,n$”) implies (10), and thus is sufficient for size-controllability of $T_{Het}$ (but not necessary, see, e.g., Example A.2).
We next explain the key observation underlying the proof of Theorem 2.2: To this end, define the (possibly empty) set of indices
$\mathcal{I}_{\#}=\left\{i:1\leq i\leq n,~R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}=0\right\},$
where $x_{i\cdot}$ denotes the $i$-th row of $X$, and define (the span of the empty set will throughout be interpreted as $\{0\}$) the space
$\mathcal{V}_{\#}=\mathop{\mathrm{s}pan}\left(\{e_{i}(n):i\in\mathcal{I}_{\#},~e_{i}(n)\in\mathsf{B}\}\right)\subseteq\mathsf{B},$ (13)
the inclusion holding because $\mathsf{B}$ is a linear space as noted earlier (recall that $R$ is $1\times k$ dimensional in this article). [^12] We note that $\mathcal{I}_{\#}$ is a proper subset of $\{1,\ldots,n\}$ since $R\neq 0$. Recall that under Assumption 1 the test statistic $T_{Het}$ as well as $\mathsf{B}$ are invariant with respect to (w.r.t.) the group $G(\mathfrak{M}_{0})$ (i.e., the group of transformations $y\mapsto\delta(y-\mu_{0})+\mu_{0}^{\ast}$ with $\delta\in\mathbb{R}$ nonzero and $\mu_{0}$ and $\mu_{0}^{\ast}$ in $\mathfrak{M}_{0}$), see Remark C.1 in Appendix C of Pötscher and Preinerstorfer (2025). [^13] The invariance holds trivially if Assumption 1 is violated. The results in Pötscher and Preinerstorfer (2025) are based on this invariance property. The crucial observation exploited in the proof of Theorem 2.2 now is that, in the special case of testing a single restriction considered in this article, the test statistic $T_{Het}$ as well as $\mathsf{B}$ are invariant, not only w.r.t. $G(\mathfrak{M}_{0})$, but also w.r.t. addition of elements of $\mathcal{V}_{\#}$. This additional invariance property involving $\mathcal{V}_{\#}$, paired with a careful application of the general theory for size-controlling critical values in Pötscher and Preinerstorfer (2018), then allows us to deduce the refined statement in Theorem 2.2. It turns out fortunate that the general theory in Pötscher and Preinerstorfer (2018) explicitly allows one to incorporate additional invariance properties beyond $G(\mathfrak{M}_{0})$. For details and proofs the reader is referred to Appendices A and B.
Finally, we remark that Theorem 2.2 is deduced from Theorem B.2 in Appendix B, which is a more general statement that also allows for heteroskedasticity models other than $\mathfrak{C}_{Het}$ (and which are defined in (14) below).
Remark 2.2: (Extensions to non-Gaussian errors) (i) All the theorems in this article continue to hold as they stand, if the disturbance vector $\mathbf{U}$ follows an elliptically symmetric distribution that has no atom at the origin; more precisely, $\mathbf{U}$ is assumed to be distributed as $\sigma\Sigma^{1/2}\mathbf{z}$, where $\mathbf{z}$ has a spherically symmetric distribution on $\mathbb{R}^{n}$ that has no atom at the origin, and where $\sigma$ and $\Sigma$ are as in Section 2.1. This is so, since the size under Gaussianity is the same as the size under the elliptical symmetry assumption. In particular, the smallest size-controlling critical values under the elliptical symmetry assumption coincide with the smallest size-controlling critical values under Gaussianity, and thus can be computed from the algorithms relying on Gaussianity described in Pötscher and Preinerstorfer (2025). See Appendix E.1 of Pötscher and Preinerstorfer (2018) and Section 7.1(i) of Pötscher and Preinerstorfer (2025) for more details. The same is actually true for a wider class of distribution for $\mathbf{U}$, namely where $\mathbf{z}$ has a distribution in the class $Z_{ua}$ defined in Appendix E.1 of Pötscher and Preinerstorfer (2018).
(ii) All the theorems in this article except for Theorem B.2 in Appendix B (i.e., all theorems using the heteroskedasticity model $\mathfrak{C}_{Het}$) continue to hold as they stand, if it is assumed that the disturbance vector $\mathbf{U}$ follows a distribution from the semiparametric model defined in Section 7.1(iv) in Pötscher and Preinerstorfer (2025) (a model that contains inter alia all distributions corresponding to i.i.d. samples of scale-mixtures of normals). Again, this is so since the size under Gaussianity is the same as the size under this semiparametric model. In particular, the smallest size-controlling critical values under this semiparametric model coincide with the smallest size-controlling critical values under Gaussianity, and thus can be computed from the algorithms relying on Gaussianity described in Pötscher and Preinerstorfer (2025). See Section 7.1(iv) in Pötscher and Preinerstorfer (2025) and note that the Gaussian model is a submodel of the semiparametric model considered there.
(iii) Furthermore, as discussed in detail in Appendix E.2 of Pötscher and Preinerstorfer (2018), any condition sufficient for size controllability under Gaussianity of the disturbance vector $\mathbf{U}$ also implies size controllability for large classes of distributions for $\mathbf{U}$ that satisfy appropriate domination conditions; however, the corresponding size-controlling critical values may then differ from the size-controlling critical values that apply under Gaussianity.
3 Conclusion
In the case of testing a single restriction, we have shown that the sufficient condition for size controllability of heteroskedasticity robust test statistics in Pötscher and Preinerstorfer (2025) can be replaced by a weaker sufficient condition that is also necessary. This allows one – in the case of testing a single restriction – to resolve the question of existence of (finite) size-controlling critical values in all cases, including those that remain inconclusive under the results in Pötscher and Preinerstorfer (2025).
We finally remark that the algorithms designed to compute size-controlling critical values as discussed in Section 10 and Appendix E of Pötscher and Preinerstorfer (2025) can be used as they stand also in situations where (a single restriction is tested and) size controllability has been verified through checking condition (10) and appealing to Theorem 2.2, but where (7) does not hold. This is so since the discussion of the before mentioned algorithms in Pötscher and Preinerstorfer (2025) only requires existence of a (finite) size-controlling critical value, but does not depend on the way this existence is verified.
Appendix A Auxiliary results
As a point of interest we note that Lemmata A.1, A.2, and A.4 below do not rely on Assumption 1. Furthermore, all the lemmata in this appendix do neither refer to the heteroskedasticity model nor to the Gaussianity assumption at all. Finally, recall from Section 2.2 that the set $\mathsf{B}$ is a linear space (as $R$ is $1\times k$ in the present article).
Lemma A.1.
The following statements hold:.
  1. 1.
    $\mathsf{B}=\mathop{\mathrm{s}pan}(X)\oplus\{\hat{u}(y):y\in\mathsf{B}\}$, the sum being orthogonal.
  2. 2.
    $\{\hat{u}(y):y\in\mathsf{B}\}$ is a linear subspace of $\mathop{\mathrm{s}pan}(e_{i}(n):i\in\mathcal{I}_{\#})$.
  3. 3.
    For every $z\in\mathop{\mathrm{s}pan}(e_{i}(n):i\in\mathcal{I}_{\#})$ we have $R\hat{\beta}(z)=0$.
  4. 4.
    If $j\in\mathcal{I}_{\#}^{c}$, then $e_{j}(n)\in\mathop{\mathrm{s}pan}(X)$ and $e_{j}(n)\in\mathsf{B}$ are equivalent.
Proof: 1. Obviously, $\{\hat{u}(y):y\in\mathsf{B}\}$ is a linear space, since $\mathsf{B}$ is so. Observe that $\hat{u}(\hat{u}(y))=\hat{u}(y)$ holds, from which it follows that $B(y)=B(\hat{u}(y))$. Consequently, $y\in\mathsf{B}$ implies $\hat{u}(y)\in\mathsf{B}$. Since $\mathsf{B}$ is invariant under addition of elements of $\mathop{\mathrm{s}pan}(X)$, we obtain $\mathsf{B}\supseteq\mathop{\mathrm{s}pan}(X)\oplus\{\hat{u}(y):y\in\mathsf{B}\}$, the sum obviously being orthogonal. For the reverse inclusion, write $y\in\mathsf{B}$ as $y=X\hat{\beta}(y)+\hat{u}(y)$, which immediately implies that $y\in\mathop{\mathrm{s}pan}(X)\oplus\{\hat{u}(y):y\in\mathsf{B}\}$.
2. Let $y\in\mathsf{B}$, i.e., $B(y)=0$, or, in other words, $R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}\hat{u}_{i}(y)=0$ for every i=1,,ni=1,...,n. It follows that $\hat{u}_{i}(y)=0$ for every $i\notin\mathcal{I}_{\#}$, from which we conclude $\hat{u}(y)\in\mathop{\mathrm{s}pan}(e_{i}(n):i\in\mathcal{I}_{\#})$.
3. With $z_{i}$ denoting the $i$-th coordinate of $z$, we have
$R\hat{\beta}(z)$ $=$ $R(X^{\prime}X)^{-1}X^{\prime}z=R(X^{\prime}X)^{-1}\mathop{\displaystyle\sum}\limits_{i=1}^{n}z_{i}x_{i\cdot}^{\prime}=\mathop{\displaystyle\sum}\limits_{i=1}^{n}z_{i}R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}$
$=$ $\mathop{\displaystyle\sum}\limits_{i\in\mathcal{I}_{\#}}z_{i}R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}+\mathop{\displaystyle\sum}\limits_{i\in\mathcal{I}_{\#}^{c}}z_{i}R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}=0,$
observing that $R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}=0$ for $i\in\mathcal{I}_{\#}$ and that $z_{i}=0$ for $i\in\mathcal{I}_{\#}^{c}$.
4. Follows from the first two claims upon noting that $j\in\mathcal{I}_{\#}^{c}$ is equivalent to $e_{j}(n)\bot\mathop{\mathrm{s}pan}(e_{i}(n):i\in\mathcal{I}_{\#})$. $\blacksquare$
Remark A.1: We discuss a few simple consequences of the preceding lemma.
(i) If $\mathcal{I}_{\#}$ is empty then $\mathsf{B}=\mathop{\mathrm{s}pan}(X)$.
(ii) If $\mathcal{I}_{\#}=\{i_{0}\}$, then $\mathsf{B}=\mathop{\mathrm{s}pan}(X)$ or $\mathsf{B}=\mathop{\mathrm{s}pan}(X)\oplus\mathop{\mathrm{s}pan}(e_{i_{0}}(n))$; the former happens if the $i_{0}$-th row of $X$ is nonzero, and the latter happens if this row is zero.
(iii) If $\mathcal{I}_{\#}$ contains more than one element, then $\mathsf{B}=\mathop{\mathrm{s}pan}(X)$ (see (iv) below) as well as $\mathsf{B}\supsetneq\mathop{\mathrm{s}pan}(X)$ (see Example A.1 below) can occur.
(iv) Suppose k=n1k=n-1 and that Assumption 1 holds. Then $\mathsf{B}=\mathop{\mathrm{s}pan}(X)$ always holds (since $\mathsf{B}$ is a linear space containing the n1n-1 dimensional subspace $\mathop{\mathrm{s}pan}(X)$ and since $\mathsf{B}$ must be a proper subspace under Assumption 1, see Lemma 3.1 in Pötscher and Preinerstorfer (2025)) regardless of whether $\mathcal{I}_{\#}$ is empty or not. [That $\mathcal{I}_{\#}$ can indeed be nonempty in this situation is shown by the example where $n=4$, $k=3$, $R=(1,1,0)^{\prime}$, and $X$ has columns $^{\prime}, (1,1,1,1)(1,-1,1,-1)^{\prime}$, and $(1,1,-1,-1)^{\prime}$. It is easy to see that $e_{i}(4)\notin\mathop{\mathrm{s}pan}(X)$ for every $i=1,\ldots,4$, and thus Assumption 1 is satisfied. The set $\mathcal{I}_{\#}$ is easily computed to be $\{2,4\}$.]
Lemma A.2.
The following statements hold:
  1. 1.
    The map $B$ and the set $\mathsf{B}$ are invariant w.r.t. addition of elements of $\mathsf{B}$. In particular, they are invariant w.r.t. addition of elements of $\mathcal{L}_{\#}:=\mathop{\mathrm{s}pan}(\mathfrak{M}_{0}^{lin}\cup\mathcal{V}_{\#})$.
  2. 2.
    $T_{Het}$ is invariant w.r.t. addition of any $z\in\mathsf{B}$ that satisfies $R\hat{\beta}(z)=0$.
  3. 3.
    $T_{Het}$ is invariant w.r.t. addition of elements of $\mathcal{L}_{\#}=\mathop{\mathrm{s}pan}(\mathfrak{M}_{0}^{lin}\cup\mathcal{V}_{\#})$.
Proof: 1. Linearity of $B:\mathbb{R}^{n}\rightarrow(\mathbb{R}^{n})^{\prime}$ together with $B(z)=0$ for every $z\in\mathsf{B}$ proves the first statement in Part 1. [The invariance claim regarding $\mathsf{B}$ also trivially follows since $\mathsf{B}$ is a linear space.] The second one then follows since, noting that $\mathsf{B}$ being a linear space, $\mathfrak{M}_{0}^{lin}\subseteq\mathop{\mathrm{s}pan}(X)\subseteq$ $\mathsf{B}$ and (13) imply $\mathcal{L}_{\#}\subseteq\mathsf{B}$.
2. First note that for $y\in\mathbb{R}^{n}$ and $z\in\mathsf{B}$ we have $\hat{\Omega}_{Het}(y+z)=\hat{\Omega}_{Het}(y)$ which follows from the easily checked representation $\hat{\Omega}_{Het}(\cdot)=B(\cdot)\mathop{\mathrm{d}iag}(d_{1},\ldots,d_{n})B^{\prime}(\cdot)$ and Part 1 of the present lemma. Second, clearly $R\hat{\beta}(y+z)-r=R\hat{\beta}(y)+R\hat{\beta}(z)-r=R\hat{\beta}(y)-r$ holds for $z$ satisfying $R\hat{\beta}(z)=0$. The claim now follows from the definition of $T_{Het}$.
[…]
Lemma A.3.
Under Assumption 1 we have $\mathop{\mathrm{d}im}(\mathcal{L}_{\#})<n-1.$
Proof: As shown in the proof of Part 1 of Lemma A.2, the relation $\mathcal{L}_{\#}\subseteq\mathsf{B}$ holds. Because $\mathsf{B}$ is a proper linear subspace of $\mathbb{R}^{n}$ under Assumption 1 (cf. Lemma 3.1 in Pötscher and Preinerstorfer (2025) and note that we have $q=1$ here), we must have $\mathop{\mathrm{d}im}(\mathcal{L}_{\#})\leq n-1$. [^14] Alternatively, $\mathop{\mathrm{d}im}(\mathcal{L}_{\#})=n$ and invariance under addition of elements of $\mathcal{L}_{\#}$ would lead to constancy of $T_{Het}$, and thus to a contradicition similar to the one arrived at in the proof in the case $\mathop{\mathrm{d}im}(\mathcal{L}_{\#})=n-1$. Assume now that $\mathcal{L}_{\#}$ has dimension n1n-1. Denote by $v\neq 0$ a vector that spans $\mathcal{L}_{\#}^{\bot}$, the orthogonal complement of $\mathcal{L}_{\#}$ in $\mathbb{R}^{n}$, and fix an arbitrary $\mu_{0}\in\mathfrak{M}_{0}$. Use the invariance property in Part 3 of Lemma A.2 to see that for every $y\notin\mathcal{L}_{\#}$ we can write
$T_{Het}(\mu_{0}+y)=T_{Het}(\mu_{0}+\Pi_{\mathcal{L}_{\#}^{\bot}}y)=T_{Het}(\mu_{0}+v),$
where we used $\Pi_{\mathcal{L}_{\#}^{\bot}}y\neq 0$ together with invariance of $T_{Het}$ w.r.t. $G(\mathfrak{M}_{0})$ (cf. Remark C.1 in Appendix C of Pötscher and Preinerstorfer (2025)) to conclude the second equality. [^15] Since $y\notin\mathcal{L}_{\#}$ we have $\Pi_{\mathcal{L}_{\#}^{\bot}}y\neq 0,$ and thus $\Pi_{\mathcal{L}_{\#}^{\bot}}y=\lambda v$ with $\lambda\neq 0$. Invariance w.r.t. the group $G(\mathfrak{M}_{0})$ then gives $T_{Het}(\mu_{0}+v)=T_{Het}(\mu_{0}+\lambda v)$. But this implies that $T_{Het}(\cdot)=T_{Het}(\mu_{0}+v)$ almost everywhere w.r.t. Lebesgue measure on $\mathbb{R}^{n}$, contradicting Part 2 of Lemma 5.16 in Pötscher and Preinerstorfer (2018) in view of Remark C.1 in Appendix C of Pötscher and Preinerstorfer (2025) and noting that Assumption 1 is being maintained. [^16] That $\mathop{\mathrm{d}im}(\mathcal{L}_{\#})=n-1$ leads to Lebesgue almost everywhere constancy has been noted in Remark 5.14(i) of Pötscher and Preinerstorfer (2018) for a large class of test statistics. We have included a proof here for the convenience of the reader. $\blacksquare$
Remark A.2: Even without Assumption 1 we always have $\mathop{\mathrm{d}im}(\mathcal{L}_{\#})<n$. To see this, note that $I_{0}(\mathcal{L}_{\#})$ is a proper subset of $\{1,\ldots,n\}$ by Part 3 of Lemma A.4 below, and thus $I_{1}(\mathcal{L}_{\#})\neq\emptyset$. But this means that $e_{i}(n)\notin\mathcal{L}_{\#}$ for at least one $i$, establishing the claim.
Lemma A.4.
The following statements hold:
  1. 1.
    $i\in\mathcal{I}_{\#}$ if and only if $\Pi_{\mathop{\mathrm{s}pan}(X)}e_{i}(n)\in\mathfrak{M}_{0}^{lin}$.
  2. 2.
    Suppose $e_{i}(n)\in\mathop{\mathrm{s}pan}(X)$. Then $i\in\mathcal{I}_{\#}$ if and only if $i\in I_{0}(\mathfrak{M}_{0}^{lin})$.
  3. 3.
    $I_{0}(\mathfrak{M}_{0}^{lin})\subseteq I_{0}(\mathcal{L}_{\#})\subseteq\mathcal{I}_{\#}$ holds, and $\mathcal{I}_{\#}$ is a proper subset of $\{1,\ldots,n\}$.
Proof: 1. Observe that
$R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}$ $=$ $R(X^{\prime}X)^{-1}X^{\prime}e_{i}(n)=R(X^{\prime}X)^{-1}X^{\prime}(\Pi_{\mathop{\mathrm{s}pan}(X)}e_{i}(n)+\Pi_{\mathop{\mathrm{s}pan}(X)^{\bot}}e_{i}(n))$
$=$ $R(X^{\prime}X)^{-1}X^{\prime}\Pi_{\mathop{\mathrm{s}pan}(X)}e_{i}(n)=R\gamma^{(i)},$
where $\gamma^{(i)}\in\mathbb{R}^{k}$ satisfies $\Pi_{\mathop{\mathrm{s}pan}(X)}e_{i}(n)=X\gamma^{(i)}$. Consequently, $i\in\mathcal{I}_{\#}$ (i.e., $R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}=0$) if and only if $R\gamma^{(i)}=0$ which is tantamount to $\Pi_{\mathop{\mathrm{s}pan}(X)}e_{i}(n)\in\mathfrak{M}_{0}^{lin}$.
2. Follows immediately from Part 1 and the definition of $I_{0}(\mathfrak{M}_{0}^{lin})$ upon noting that $\Pi_{\mathop{\mathrm{s}pan}(X)}e_{i}(n)=e_{i}(n)$ because of the assumption $e_{i}(n)\in\mathop{\mathrm{s}pan}(X)$.
3. The first inclusion is trivial since $\mathfrak{M}_{0}^{lin}\subseteq\mathcal{L}_{\#}$. To prove the second inclusion, suppose $i\in I_{0}(\mathcal{L}_{\#})$. Then $e_{i}(n)\in\mathcal{L}_{\#}$, which implies that $e_{i}(n)=v+w$ where $v\in\mathcal{V}_{\#}$ and $w\in\mathfrak{M}_{0}^{lin}$ (here we also use that $\mathcal{V}_{\#}$ and $\mathfrak{M}_{0}^{lin}$ are linear subspaces). Using the definition of $\mathcal{V}_{\#}$ we arrive at
$e_{i}(n)=\sum_{j:j\in\mathcal{I}_{\#},e_{j}(n)\in\mathsf{B}}\lambda_{j}e_{j}(n)+w.$
Taking the projection and noting that $\Pi_{\mathop{\mathrm{s}pan}(X)}w=w$ (since $w\in\mathfrak{M}_{0}^{lin}\subseteq\mathop{\mathrm{s}pan}(X)$) this gives
$\Pi_{\mathop{\mathrm{s}pan}(X)}e_{i}(n)=\sum_{j:j\in\mathcal{I}_{\#},e_{j}(n)\in\mathsf{B}}\lambda_{j}\Pi_{\mathop{\mathrm{s}pan}(X)}e_{j}(n)+w.$
The already established Part 1 shows that $\Pi_{\mathop{\mathrm{s}pan}(X)}e_{j}(n)\in\mathfrak{M}_{0}^{lin}$ for $j\in\mathcal{I}_{\#}$. Since $\mathfrak{M}_{0}^{lin}$ is a linear space we conclude that $\Pi_{\mathop{\mathrm{s}pan}(X)}e_{i}(n)$ belongs to $\mathfrak{M}_{0}^{lin}$. Again using Part 1, we arrive at $i\in\mathcal{I}_{\#}$. That $\mathcal{I}_{\#}$ is a proper subset of $\{1,\ldots,n\}$ follows since $R\neq 0$. $\blacksquare$
Remark A.3: (i) Example A.1 below and the example discussed towards the end of Remark A.1(iv) show that the first two inclusions in Part 3 of the above lemma can be strict inclusions.
(ii) Inspection of the proof shows that Lemma A.4 actually also holds if, in the notation of Pötscher and Preinerstorfer (2025), we have $q\geq 1$, i.e., if a collection of $q$ restrictions is tested simultaneously.
The subsequent examples show that condition (7) can be stronger than condition (10), another such example being Example C.1 in Appendix C.1 of Pötscher and Preinerstorfer (2025). We provide four different examples to illustrate that this can happen in a variety of different situations (e.g., independently of whether standard basis vectors belong to $\mathop{\mathrm{s}pan}(X)$ or not, etc.). We also compute the set $\mathsf{B}$ in the examples below and illustrate the results in Lemma A.1.
Example A.1: Suppose $k=2$, $n=4$, and $X$ has $(1,1,1,1)^{\prime}$ as its first column and $(1,-1,1,-1)^{\prime}$ as its second column. Define the $1\times k$ vector $R=(1,1)$. Then $\mathop{\mathrm{r}ank}(X)=k=2$ holds, and $e_{j}(4)\notin\mathop{\mathrm{s}pan}(X)$ for every $j=1,\ldots,4$, as is easily checked; in particular, Assumption 1 is thus satisfied, and $I_{1}(\mathfrak{M}_{0}^{lin})=\{1,\ldots,4\}$. Furthermore, $R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}\neq 0$ for $i=1,3$ whereas $R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}=0$ for $i=2,4$. I.e., $\mathcal{I}_{\#}=\{2,4\}$. Now, $y\in\mathsf{B}$ (i.e., $B(y)=0$) is easily seen to be equivalent to $\hat{u}_{1}(y)=\hat{u}_{3}(y)=0$, which in turn is equivalent to $y_{1}=y_{3}$. In particular, $e_{2}(4)$ and $e_{4}(4)$ belong to $\mathsf{B}$, but do not belong to $\mathop{\mathrm{s}pan}(X)$, while $e_{1}(4)$ and $e_{3}(4)$ do not belong to $\mathsf{B}$. The space $\{\hat{u}(y):y\in\mathsf{B}\}$ in the orthogonal sum representation $\mathsf{B}=\mathop{\mathrm{s}pan}(X)\oplus\{\hat{u}(y):y\in\mathsf{B}\}$ is here given by $\mathop{\mathrm{s}pan}((0,1,0,-1)^{\prime})$ as is not difficult to see. Note that, while $e_{2}(4)$ and $e_{4}(4)$ belong to $\mathsf{B}$ (and trivially also to $\mathop{\mathrm{s}pan}(e_{i}(4):i\in\mathcal{I}_{\#})$), they are not orthogonal to $\mathop{\mathrm{s}pan}(X)$, and do not belong to $\mathop{\mathrm{s}pan}((0,1,0,-1)^{\prime})$ (which is a subset of $\mathop{\mathrm{s}pan}(e_{i}(4):i\in\mathcal{I}_{\#})$). Furthermore, since $I_{1}(\mathfrak{M}_{0}^{lin})=\{1,\ldots,4\}$, condition (10) is satisfied, while condition (7) is not. Theorem 2.1 does not allow one to draw a conclusion about size-controllability of $T_{Het}$ in this example, while Theorem 2.2 shows that $T_{Het}$ is size-controllable.
Example A.2: Suppose $k=3$, $n=5$, and $X$ has $(1,1,1,1,0)^{\prime}$ as its first column, $(1,-1,1,-1,0)^{\prime}$ as its second column, and $(0,0,0,0,2)^{\prime}$ as its last column. Define the $1\times k$ vector $R=(1,1,r_{3})$. Then $\mathop{\mathrm{r}ank}(X)=k=3$ holds, and $e_{j}(5)\notin\mathop{\mathrm{s}pan}(X)$ for every $j=1,\ldots,4$, but $e_{5}(5)\in\mathop{\mathrm{s}pan}(X)$. Assumption 1 is satisfied as can be easily checked. Furthermore, $R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}\neq 0$ for $i=1,3$, whereas $R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}=0$ for $i=2,4$; and $R(X^{\prime}X)^{-1}x_{5\cdot}^{\prime}=r_{3}/2$. Hence, $\mathcal{I}_{\#}=\{2,4\}$ in case $r_{3}\neq 0$, and $\mathcal{I}_{\#}=\{2,4,5\}$ otherwise. Now, $y\in\mathsf{B}$ (i.e., $B(y)=0$) is easily seen to be equivalent to $\hat{u}_{1}(y)=\hat{u}_{3}(y)=0$, which in turn is equivalent to $y_{1}=y_{3}$. In particular, $e_{2}(5)$ and $e_{4}(5)$ belong to $\mathsf{B}$, but do not belong to $\mathop{\mathrm{s}pan}(X)$, while $e_{5}(5)\in\mathop{\mathrm{s}pan}(X)\subseteq\mathsf{B}$; and $e_{1}(5)$ and $e_{3}(5)$ do not belong to $\mathsf{B}$. The space $\{\hat{u}(y):y\in\mathsf{B}\}$ in the orthogonal sum representation $\mathsf{B}=\mathop{\mathrm{s}pan}(X)\oplus\{\hat{u}(y):y\in\mathsf{B}\}$ is here given by $\mathop{\mathrm{s}pan}((0,1,0,-1,0)^{\prime})$ as is not difficult to see. Note that, while $e_{2}(5)$ and $e_{4}(5)$ belong to $\mathsf{B}$ (and trivially also to $\mathop{\mathrm{s}pan}(e_{i}(5):i\in\mathcal{I}_{\#})$), they are not orthogonal to $\mathop{\mathrm{s}pan}(X)$, and do not belong to $\mathop{\mathrm{s}pan}((0,1,0,-1,0)^{\prime})$ (which is a subset of $\mathop{\mathrm{s}pan}(e_{i}(5):i\in\mathcal{I}_{\#})$). Note that $I_{1}(\mathfrak{M}_{0}^{lin})=\{1,\ldots,4\}$ in case $r_{3}=0$, while $I_{1}(\mathfrak{M}_{0}^{lin})=\{1,\ldots,5\}$ otherwise. In particular, in case $r_{3}=0$, condition (10) is satisfied, while condition (7) is not; hence, in this case Theorem 2.1 does not allow one to draw a conclusion about size-controllability of $T_{Het}$, while Theorem 2.2 shows that $T_{Het}$ is size-controllable. In case $r_{3}\neq 0$, both conditions (7) and (10) are violated, and both theorems show that the test based on $T_{Het}$ has size $1$ regardless of the choice of critical value.
Example A.3: Suppose $k=2$, $n=5$, and $X$ has $(1,1,1,1,0)^{\prime}$ as its first column and $(1,-1,1,-1,0)^{\prime}$ as its second column. Define the $1\times k$ vector $R=(1,0)$. Then $\mathop{\mathrm{r}ank}(X)=k=2$ holds, and $e_{j}(5)\notin\mathop{\mathrm{s}pan}(X)$ for every $j=1,\ldots,5$, as is easily checked; in particular, Assumption 1 is thus satisfied, and $I_{1}(\mathfrak{M}_{0}^{lin})=\{1,\ldots,5\}$. Furthermore, $R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}\neq 0$ for $i=1,\ldots,4$ whereas $R(X^{\prime}X)^{-1}x_{5\cdot}^{\prime}=0$. I.e., $\mathcal{I}_{\#}=\{5\}$. Now, $y\in\mathsf{B}$ (i.e., $B(y)=0$) is easily seen to be equivalent to $\hat{u}_{1}(y)=\hat{u}_{2}(y)=\hat{u}_{3}(y)=\hat{u}_{4}(y)=0$, which in turn is equivalent to $y_{1}=y_{3}$ and $y_{2}=y_{4}$. In particular, $e_{5}(5)$ belongs to $\mathsf{B}$, but does not belong to $\mathop{\mathrm{s}pan}(X)$, in fact is orthogonal to $\mathop{\mathrm{s}pan}(X)$, while $e_{j}(5)\notin\mathsf{B}$ for $j=1,\ldots,4$. The space $\{\hat{u}(y):y\in\mathsf{B}\}$ in the orthogonal sum representation $\mathsf{B}=\mathop{\mathrm{s}pan}(X)\oplus\{\hat{u}(y):y\in\mathsf{B}\}$ is here given by $\mathop{\mathrm{s}pan}(e_{5}(5))$ as is not difficult to see. Furthermore, in this example condition (10) is satisfied, while condition (7) is not. Theorem 2.1 does not allow one to draw a conclusion about size-controllability of $T_{Het}$ in this example, while Theorem 2.2 shows that $T_{Het}$ is size-controllable.
Example A.4: Suppose $k=3$, $n=6$, and $X$ has $(1,1,1,1,0,0)^{\prime}$ as its first column, $(1,-1,1,-1,0,0)^{\prime}$ as its second column, and $(0,0,0,0,0,2)^{\prime}$ as its third column. Define the $1\times k$ vector $R=(1,0,r_{3})$. Then $\mathop{\mathrm{r}ank}(X)=k=3$ holds, and $e_{j}(6)\notin\mathop{\mathrm{s}pan}(X)$ for every $j=1,\ldots,5$, but $e_{6}(6)\in\mathop{\mathrm{s}pan}(X)$. Assumption 1 is satisfied as can be easily checked. Furthermore, $R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}\neq 0$ for $i=1,\ldots,4$ whereas $R(X^{\prime}X)^{-1}x_{5\cdot}^{\prime}=0$ and $R(X^{\prime}X)^{-1}x_{6\cdot}^{\prime}=r_{3}/2$. Hence, $\mathcal{I}_{\#}=\{5\}$ in case $r_{3}\neq 0,$ and $\mathcal{I}_{\#}=\{5,6\}$ otherwise. Now, $y\in\mathsf{B}$ (i.e., $B(y)=0$) is easily seen to be equivalent to $\hat{u}_{1}(y)=\hat{u}_{2}(y)=\hat{u}_{3}(y)=\hat{u}_{4}(y)=0$, which in turn is equivalent to $y_{1}=y_{3}$ and $y_{2}=y_{4}$. In particular, $e_{5}(6)$ belongs to $\mathsf{B}$, but does not belong to $\mathop{\mathrm{s}pan}(X)$, in fact is orthogonal to $\mathop{\mathrm{s}pan}(X)$, while $e_{6}(6)\in\mathop{\mathrm{s}pan}(X)\subseteq\mathsf{B}$; and $e_{j}(6)\notin\mathsf{B}$ for $j=1,\ldots,4$. The space $\{\hat{u}(y):y\in\mathsf{B}\}$ in the orthogonal sum representation $\mathsf{B}=\mathop{\mathrm{s}pan}(X)\oplus\{\hat{u}(y):y\in\mathsf{B}\}$ is here given by $\mathop{\mathrm{s}pan}(e_{5}(6))$ as is not difficult to see. Note that $I_{1}(\mathfrak{M}_{0}^{lin})=\{1,\ldots,5\}$ in case $r_{3}=0$, while $I_{1}(\mathfrak{M}_{0}^{lin})=\{1,\ldots,6\}$ otherwise. In particular, in case $r_{3}=0$, condition (10) is satisfied, while condition (7) is not; hence, in this case Theorem 2.1 does not allow one to draw a conclusion about size-controllability of $T_{Het}$, while Theorem 2.2 shows that $T_{Het}$ is size-controllable. In case $r_{3}\neq 0$, both conditions (7) and (10) are violated, and both theorems show that the test based on $T_{Het}$ has size $1$ regardless of the choice of critical value.
Remark A.4: Many more examples can be generated from Examples A.1-A.4 via the transformation $X^{\ast}=XA$ and $R^{\ast}=RA$ where $A$ is a nonsingular $k\times k$ matrix. These new examples exhibit the same features as Examples A.1-A.4, respectively. In particular, one can generate examples that have $R^{\ast}=(1,0\ldots,0)$.
Appendix B Proof of Theorem 2.2
To prove Theorem 2.2 we follow the strategy used to establish Theorem 5.1 in Pötscher and Preinerstorfer (2025) and first provide a result for a class of heteroskedasticity models that includes $\mathfrak{C}_{Het}$ as a special case, and which is of some independent interest. The heteroskedasticity models we consider here are defined as follows (cf. Appendix A of Pötscher and Preinerstorfer (2025) for more discussion): Let $m\in\mathbb{N}$, and let $n_{j}\in\mathbb{N}$ for $j=1,\ldots,m$ satisfy $\sum_{j=1}^{m}n_{j}=n$. Set $n_{j}^{+}=\sum_{l=1}^{j}n_{l}$ and define
$\mathfrak{C}_{(n_{1},\ldots,n_{m})}=\left\{\mathop{\mathrm{d}iag}(\tau_{1}^{2},\ldots,\tau_{n}^{2})\in\mathfrak{C}_{Het}:\tau_{n_{j-1}^{+}+1}^{2}=\ldots=\tau_{n_{j}^{+}}^{2}\text{ for }j=1,\ldots,m\right\}$ (14)
with the convention that $n_{0}^{+}=0$. In the special case where $m=n$ and $n_{1}=n_{2}=...=n_{m}=1$ we have $\mathfrak{C}_{(n_{1},\ldots,n_{m})}=\mathfrak{C}_{Het}$. We use $\lambda_{\mathbb{R}^{n}}$ to denote Lebesgue measure on $\mathbb{R}^{n}$, and $\lambda_{\mathcal{A}}$ to denote Lebesgue measure on a (nonempty) affine space $\mathcal{A}$ (but viewed as a measure on the Borel-sets of $\mathbb{R}^{n}$), with zero-dimensional Lebesgue measure interpreted as point mass. We start with a lemma and note that it does not make use of Assumption 1. Recall that by definition $\mathcal{L}_{\#}=\mathop{\mathrm{s}pan}(\mathfrak{M}_{0}^{lin}\cup\mathcal{V}_{\#})$, and that we only consider testing a single restriction in the present article.
Lemma B.1.
Let $m\in\mathbb{N}$, and let $n_{j}\in\mathbb{N}$ for $j=1,\ldots,m$ satisfy $\sum_{j=1}^{m}n_{j}=n$. Then:
(a) The condition
$\mathop{\mathrm{s}pan}\left(\left\{e_{i}(n):i\in(n_{j-1}^{+},n_{j}^{+}]\right\}\right)$ $\nsubseteqq\mathsf{B}\text{ }$
$\text{\ for every }j$ $=1,\ldots,m\text{ with }(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})\neq\emptyset$ (15)
is equivalent to the condition
$\mathop{\mathrm{s}pan}\left(\left\{e_{i}(n):i\in(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#}\right\}\right)$ $\nsubseteqq\mathop{\mathrm{s}pan}(X)$
$\text{ for every }j$ $=1,\ldots,m\text{ with }\emptyset\neq(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})\subseteq\mathcal{I}_{\#}^{c}.$ (16)
[It is understood here, that condition (16) is satisfied if no $j$ with $\emptyset\neq(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})\subseteq\mathcal{I}_{\#}^{c}$ exists.]
(b) In the special case where $m=n$ and $n_{1}=n_{2}=...=n_{m}=1$, (16) (as well as (15)) is equivalent to (10).
Proof: (a) Recall from the proof of Part 1 of Lemma A.2 that $\mathcal{L}_{\#}=\mathop{\mathrm{s}pan}(\mathfrak{M}_{0}^{lin}\cup\mathcal{V}_{\#})\subseteqq\mathsf{B}$. Therefore, $e_{i}(n)\notin\mathsf{B}$ is possible only if $i\in I_{1}(\mathcal{L}_{\#})$. Hence, in view of invariance of $\mathsf{B}$ w.r.t. addition of elements of $\mathsf{B}$ (Lemma A.2), the condition in (15) is equivalent to
$\mathop{\mathrm{s}pan}\left(\left\{e_{i}(n):i\in(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})\right\}\right)$ $\nsubseteqq$ $\mathsf{B}$
$\text{\ for every }j$ $=$ $1,\ldots,m\text{ with }(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})\neq\emptyset.$ (17)
For $i\in\mathcal{I}_{\#}$ the condition $e_{i}(n)\in\mathsf{B}$ implies $e_{i}(n)\in\mathcal{V}_{\#}\subseteq\mathcal{L}_{\#}$, so that $i\notin I_{1}(\mathcal{L}_{\#})$. In other words, $i\in I_{1}(\mathcal{L}_{\#})\cap\mathcal{I}_{\#}$ implies $e_{i}(n)\notin\mathsf{B}$. This shows that for any $j$ with the property that $(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})$ contains an element $i\in\mathcal{I}_{\#}$, the non-inclusion relation in (17) is automatically satisfied. Hence, (17) is equivalent to
$\mathop{\mathrm{s}pan}\left(\left\{e_{i}(n):i\in(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})\right\}\right)$ $\nsubseteqq$ $\mathsf{B}$ (18)
$\text{for every }j$ $=$ $1,\ldots,m\text{ with }\emptyset\neq(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})\subseteq\mathcal{I}_{\#}^{c}$
with the understanding that this condition is satisfied if no $j$ with $\emptyset\neq(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})\subseteq\mathcal{I}_{\#}^{c}$ exists. Since $\mathsf{B}$ as well as $\mathop{\mathrm{s}pan}(X)$ are a linear spaces, Part 4 of Lemma A.1 shows that (18) is equivalent to the statement in (16).
(b) In the special case considered here (16) simplifies to
$e_{i}(n)\notin\mathop{\mathrm{s}pan}(X)\quad\text{ \ for every }i\in I_{1}(\mathcal{L}_{\#})\cap\mathcal{I}_{\#}^{c}$ (19)
with the understanding as in (16) that this condition is satisfied if $I_{1}(\mathcal{L}_{\#})\cap\mathcal{I}_{\#}^{c}$ is empty. Since $I_{1}(\mathcal{L}_{\#})\cap\mathcal{I}_{\#}^{c}=\mathcal{I}_{\#}^{c}\neq\emptyset$ by Part 3 of Lemma A.4, the index set in (19) is actually nonempty, and furthermore (19) is equivalent to
$e_{i}(n)\notin\mathop{\mathrm{s}pan}(X)\quad\text{ \ for every }i\in\mathcal{I}_{\#}^{c}.$ (20)
Because of $\mathcal{I}_{\#}^{c}\subseteq I_{1}(\mathfrak{M}_{0}^{lin})$ (Lemma A.4), the statement in (20) is implied by that in (10). To show that (20) implies (10), suppose (10) is violated, i.e., there exists an $i\in I_{1}(\mathfrak{M}_{0}^{lin})$ such that $e_{i}(n)\in\mathop{\mathrm{s}pan}(X)$. It then follows that $R\hat{\beta}(e_{i}(n))\neq 0$ must hold. Since $R\hat{\beta}(e_{i}(n))=R(X^{\prime}X)^{-1}x_{i\cdot}^{\prime}$, we conclude $i\in\mathcal{I}_{\#}^{c}$. Hence, also (20) must be violated, a contradiction. $\blacksquare$
Parts 1-2 of the following statement provide – in the context of testing a single restriction – a version of Theorem A.1(b) and the corresponding part of Theorem A.1(c) in Pötscher and Preinerstorfer (2025), while Part 3 corresponds to the generalization of Proposition 5.5(b) mentioned after Theorem A.1 in Pötscher and Preinerstorfer (2025). Part 4 of the subsequent theorem is a version of Proposition A.2(b) in Pötscher and Preinerstorfer (2025), and together with Part 1 shows that under Assumption 1 the condition in (15), or equivalently (16), is necessary and sufficient for the existence of a (finite) critical value that controls the size of $T_{Het}$ over the heteroskedasticity model $\mathfrak{C}_{(n_{1},\ldots,n_{m})}$ when testing
$H_{0}:\mu\in\mathfrak{M}_{0},\ 0<\sigma^{2}<\infty,\ \Sigma\in\mathfrak{C}_{(n_{1},\ldots,n_{m})}\quad\text{ vs. }\quad H_{1}:\mu\in\mathfrak{M}_{1},\ 0<\sigma^{2}<\infty,\ \Sigma\in\mathfrak{C}_{(n_{1},\ldots,n_{m})}.$
Theorem B.2.
Let $m\in\mathbb{N}$, let $n_{j}\in\mathbb{N}$ for $j=1,\ldots,m$ satisfy $\sum_{j=1}^{m}n_{j}=n$, and suppose Assumption 1 is satisfied. Then the following statements hold:
  1. 1.
    For every $0<\alpha<1$ there exists a real number $C(\alpha)$ such that
    $\sup_{\mu_{0}\in\mathfrak{M}_{0}}\sup_{0<\sigma^{2}<\infty}\sup_{\Sigma\in\mathfrak{C}_{(n_{1},\ldots,n_{m})}}P_{\mu_{0},\sigma^{2}\Sigma}(T_{Het}\geq C(\alpha))\leq\alpha$ (21)
    holds, provided that (15) (or equivalently (16)) holds. Furthermore, under condition (15) (or equivalently (16)), even equality can be achieved in (21) by a proper choice of $C(\alpha)$, provided $\alpha\in(0,\alpha^{\ast}]\cap(0,1)$ holds, where
    $\alpha^{\ast}=\sup_{C\in(C^{\ast},\infty)}\sup_{\Sigma\in\mathfrak{C}_{(n_{1},\ldots,n_{m})}}P_{\mu_{0},\Sigma}(T_{Het}\geq C)$
    is positive and where $C^{\ast}$ is defined as in Lemma 5.11 of Pötscher and Preinerstorfer (2018) with $\mathfrak{C}=\mathfrak{C}_{(n_{1},\ldots,n_{m})}$, $T=T_{Het}$, $N^{{\dagger}}=\mathsf{B}$, $\mathcal{L}=\mathcal{L}_{\#}$, and $q=1$ (with neither $\alpha^{\ast}$ nor $C^{\ast}$ depending on the choice of $\mu_{0}\in\mathfrak{M}_{0}$).
  2. 2.
    Suppose (15) (or equivalently (16)) is satisfied. Then a smallest critical value, denoted by $C_{\Diamond}(\alpha)$, satisfying (21) exists for every $0<\alpha<1$. And $C_{\Diamond}(\alpha)$ is also the smallest among the critical values leading to equality in (21) whenever such critical values exist. [^17] The dependence of $C_{\Diamond}(\alpha)$ on the heteroskedasticity model is not shown in the notation, In particular, $C_{\Diamond}(\alpha)$ in the current theorem is not necessarily the same as $C_{\Diamond}(\alpha)$ in the other theorems.
  3. 3.
    Suppose (15) (or equivalently (16)) is satisfied. Then any $C(\alpha)$ satisfying (21) necessarily has to satisfy $C(\alpha)\geq C^{\ast}$. In fact, for any $C<C^{\ast}$ we have $\sup_{\Sigma\in\mathfrak{C}_{(n_{1},\ldots,n_{m})}}P_{\mu_{0},\sigma^{2}\Sigma}(T_{Het}\geq C)=1$ for every $\mu_{0}\in\mathfrak{M}_{0}$ and every $\sigma^{2}\in(0,\infty)$.
  4. 4.
    If (15) (or equivalently (16)) is violated, then $\sup_{\Sigma\in\mathfrak{C}_{(n_{1},\ldots,n_{m})}}P_{\mu_{0},\sigma^{2}\Sigma}(T_{Het}\geq C)=1$ for every choice of critical value $C$, every $\mu_{0}\in\mathfrak{M}_{0}$, and every $\sigma^{2}\in(0,\infty)$ (implying that size equals $1$ for every CC). [^18] Cf. Footnote 8.
The following proof adapts the proof of Theorem A.1 in Pötscher and Preinerstorfer (2025).
Proof of Theorem B.2: We first prove Part 1. We apply Part A of Proposition 5.12 of Pötscher and Preinerstorfer (2018) with $\mathfrak{C}=\mathfrak{C}_{(n_{1},\ldots,n_{m})}$, $T=T_{Het}$, $\mathcal{L}=\mathcal{L}_{\#}$, and $\mathcal{V}=\mathcal{V}_{\#}$ (and $q=1$). First, note that $\dim(\mathcal{L}_{\#})<n-1<n$ because of Lemma A.3. Second, under Assumption 1, $T_{Het}$ is a non-sphericity corrected F-type test statistic with $N^{\ast}=\mathsf{B}$, which is a closed $\lambda_{\mathbb{R}^{n}}$-null set (see Remarks 3.2 and C.1 as well as Lemma 3.1 in Pötscher and Preinerstorfer (2025)); in particular, $T_{Het}$ as well as $\mathsf{B}$ are invariant w.r.t. the group $G(\mathfrak{M}_{0})$. Furthermore, $T_{Het}$ as well as $\mathsf{B}$ are invariant w.r.t. addition of elements of $\mathcal{V}_{\#}$ by Lemma A.2. Hence, the general assumptions on $T=T_{Het}$, on $N^{{\dagger}}=N^{\ast}=\mathsf{B}$, on $\mathcal{V}=\mathcal{V}_{\#}$, as well as on $\mathcal{L}=\mathcal{L}_{\#}$ in Proposition 5.12 of Pötscher and Preinerstorfer (2018) are satisfied in view of Part 1 of Lemma 5.16 in the same reference.
Next, observe that condition (15) is equivalent to
$\mathop{\mathrm{s}pan}\left(\left\{\Pi_{\mathcal{L}_{\#}^{\bot}}e_{i}(n):i\in(n_{j-1}^{+},n_{j}^{+}]\right\}\right)\nsubseteqq\mathsf{B}$
for every $j=1,\ldots,m$, such that $(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})\neq\emptyset$, since $\Pi_{\mathcal{L}_{\#}^{\bot}}e_{i}(n)$ and $e_{i}(n)$ differ only by an element of $\mathcal{L}_{\#}$ and since $\mathsf{B}+\mathcal{L}_{\#}=\mathsf{B}$ (which follows from Part 1 of Lemma A.2). In view of Proposition B.2 in Appendix B of Pötscher and Preinerstorfer (2025), this implies that any $\mathcal{S}\in\mathbb{J}(\mathcal{L}_{\#},\mathfrak{C}_{(n_{1},\ldots,n_{m})})$ is not contained in $\mathsf{B}$, and thus not in $N^{{\dagger}}$. Using $\mathfrak{M}_{0}\subseteq\mathop{\mathrm{s}pan}(X)$ and $\mathsf{B}+\mathop{\mathrm{s}pan}(X)=\mathsf{B}$ (by Lemma 3.1(e) in Pötscher and Preinerstorfer (2025)), it follows that $\mu_{0}+\mathcal{S}\nsubseteqq\mathsf{B}=N^{{\dagger}}$ for every $\mu_{0}\in\mathfrak{M}_{0}$. Since $\mu_{0}+\mathcal{S}$ is a (nonempty) affine space and $N^{{\dagger}}=\mathsf{B}$ is a linear space (recall that $R$ is $1\times k$), we may conclude (cf. Corollary 5.6 in Pötscher and Preinerstorfer (2018) and its proof) that $\lambda_{\mu_{0}+\mathcal{S}}(N^{{\dagger}})=0$ for every $\mathcal{S}\in\mathbb{J}(\mathcal{L}_{\#},\mathfrak{C}_{(n_{1},\ldots,n_{m})})$ and every $\mu_{0}\in\mathfrak{M}_{0}$. This completes the verification of the assumptions of Proposition 5.12 in Pötscher and Preinerstorfer (2018) that are not specific to Part A (or Part B) of this proposition.
We next verify the assumptions specific to Part A of this proposition: Assumption (a) is satisfied (even for every $C\in\mathbb{R}$) as a consequence of Part 2 of Lemma 5.16 in Pötscher and Preinerstorfer (2018) and of Remark C.1(i) in Appendix C of Pötscher and Preinerstorfer (2025). And Assumption (b) in Part A follows from Lemma 5.19 of Pötscher and Preinerstorfer (2018), since $T_{Het}$ results as a special case of the test statistics $T_{GQ}$ defined in Section 3.4 of Pötscher and Preinerstorfer (2018) upon choosing $\mathcal{W}_{n}^{\ast}=n^{-1}\mathop{\mathrm{d}iag}(d_{1},\ldots,d_{n})$. Part A of Proposition 5.12 of Pötscher and Preinerstorfer (2018) now immediately delivers claim (21), since $C^{\ast}<\infty$ as noted in that proposition. That $C^{\ast}$ and $\alpha^{\ast}$ do not depend on the choice of $\mu_{0}\in\mathfrak{M}_{0}$ is an immediate consequence of $G(\mathfrak{M}_{0})$-invariance of $T_{Het}$ (cf. Remark 3.2 in Pötscher and Preinerstorfer (2025)). Also note that $\alpha^{\ast}$ as defined in the theorem coincides with $\alpha^{\ast}$ as defined in Proposition 5.12 of Pötscher and Preinerstorfer (2018) in view of $G(\mathfrak{M}_{0})$-invariance of $T_{Het}$. Positivity of $\alpha^{\ast}$ then follows from Part 5 of Lemma 5.15 in Preinerstorfer and Pötscher (2016) in view of Remark C.1(i) in Appendix C of Pötscher and Preinerstorfer (2025), noting that $\lambda_{\mathbb{R}^{n}}$ and $P_{\mu_{0},\Sigma}$ are equivalent measures (since $\Sigma\in\mathfrak{C}_{Het}$ is positive definite); cf. Remark 5.13(vi) in Pötscher and Preinerstorfer (2018). In case $\alpha<\alpha^{\ast}$, the remaining claim in Part 1 of the present theorem, namely that equality can be achieved in (21), follows from the definition of $C^{\ast}$ in Lemma 5.11 of Pötscher and Preinerstorfer (2018) and from Part A.2 of Proposition 5.12 of Pötscher and Preinerstorfer (2018) (and the observation immediately following that proposition allowing one to drop the suprema w.r.t. $\mu_{0}$ and $\sigma^{2}$, and to set $\sigma^{2}=1$); in case $\alpha=\alpha^{\ast}<1$, it follows from Remarks 5.13(i),(ii) in Pötscher and Preinerstorfer (2018) using Lemma 5.16 in the same reference.
The claim in Part 2 follows from Remark 5.10 and Lemma 5.16 in Pötscher and Preinerstorfer (2018) combined with Remark C.1(i) in Appendix C of Pötscher and Preinerstorfer (2025); cf. also Appendix A.3 in Pötscher and Preinerstorfer (2025).
Part 3 follows from Part A.1 of Proposition 5.12 of Pötscher and Preinerstorfer (2018) and the sentence following this proposition. Note that the assumptions of this proposition have been verified in the proof of Part 1 above.
Part 4 follows from Part 3 of Corollary 5.17 in Preinerstorfer and Pötscher (2016): As shown in Remark C.1 in Appendix C of Pötscher and Preinerstorfer (2025), $T_{Het}$ satisfies the assumptions of this corollary (with $\check{\beta}=\hat{\beta}$, $\check{\Omega}=\hat{\Omega}_{Het}$, $N=\emptyset$, and $N^{\ast}=\mathsf{B}$). Suppose that (16) is violated and set $\mathcal{Z}=\mathop{\mathrm{s}pan}(\{e_{i}(n):i\in(n_{j-1}^{+},n_{j}^{+}]\})$, where $j$ is such that $\emptyset\neq(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})\subseteq\mathcal{I}_{\#}^{c}$ and
$\mathop{\mathrm{s}pan}\left(\left\{e_{i}(n):i\in(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})\right\}\right)\subseteq\mathop{\mathrm{s}pan}(X).$ (22)
Since $e_{i}(n)\in\mathcal{L}_{\#}$ for every $i\in I_{0}(\mathcal{L}_{\#})$, it hence follows from (22) that $\mathcal{Z}\subseteq\mathop{\mathrm{s}pan}(\mathop{\mathrm{s}pan}(X)\cup\mathcal{L}_{\#})\subseteq\mathsf{B}$, recalling that $\mathop{\mathrm{s}pan}(X)\subseteq\mathsf{B}$, that $\mathcal{L}_{\#}\subseteq\mathsf{B}$ (cf. the proof of Part 1 of Lemma A.2), and that $\mathsf{B}$ is a linear space (recall that $R$ is $1\times k$). Note that $\mathcal{Z}$ is not contained in $\mathfrak{M}_{0}^{lin}$ because $\emptyset\neq(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})$ but $\mathfrak{M}_{0}^{lin}\subseteq\mathcal{L}_{\#}$. Observe that $\mathcal{Z}$ is a concentration space of $\mathfrak{C}_{(n_{1},\ldots,n_{m})}$ in view of Remark B.4 in Appendix B of Pötscher and Preinerstorfer (2025) (note that $\mathop{\mathrm{c}ard}((n_{j-1}^{+},n_{j}^{+}])<n$ must hold in view of $\mathcal{Z}\subseteq\mathsf{B}$ and $\mathsf{B}$ being a proper subspace of $\mathbb{R}^{n}$ by Lemma 3.1 in Pötscher and Preinerstorfer (2025) in conjunction with Assumption 1, while $0<\mathop{\mathrm{c}ard}((n_{j-1}^{+},n_{j}^{+}])$ is obvious). The nonnegative definiteness assumption on $\check{\Omega}=\hat{\Omega}_{Het}$ in Part 3 of Corollary 5.17 in Preinerstorfer and Pötscher (2016) is satisfied (cf. Lemma 3.1 in Pötscher and Preinerstorfer (2025)). Obviously $\check{\Omega}(z)=0$ holds for every $z\in\mathcal{Z}$ as a consequence of Part (b) of Lemma 3.1 in Pötscher and Preinerstorfer (2025) since $\mathcal{Z}\subseteq\mathsf{B}$ (as just shown) and since $\check{\Omega}(z)$ is $1\times 1$. It remains to establish that $R\hat{\beta}(z)\neq 0$ holds $\lambda_{\mathcal{Z}}$-everywhere: we recall that $\emptyset\neq(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})\subseteq\mathcal{I}_{\#}^{c}$ and pick an element $i$, say, of $(n_{j-1}^{+},n_{j}^{+}]\cap I_{1}(\mathcal{L}_{\#})$. Then $e_{i}(n)\in\mathcal{Z}$ and $i\in\mathcal{I}_{\#}^{c}$, and from the definition of $\mathcal{I}_{\#}^{c}$ we conclude that $R\hat{\beta}(e_{i}(n))\neq 0$. It follows that the linear space $\mathcal{Z}$ is not a subspace of the kernel of $R\hat{\beta}$ so that $R\hat{\beta}(z)\neq 0$ holds $\lambda_{\mathcal{Z}}$-everywhere. Part 3 of Corollary 5.17 in Preinerstorfer and Pötscher (2016) then proves the claim for $C>0$. A fortiori it then also holds for all real $C$. $\blacksquare$
We are now ready to prove Theorem 2.2. The proof follows the structure of the proof of Theorem 5.1 in Pötscher and Preinerstorfer (2025).
[…]
$C^{\ast}\geq\max\left\{T_{Het}(\mu_{0}+\Pi_{\mathcal{L}_{\#}^{\bot}}e_{i}(n)):i\in I_{1}(\mathcal{L}_{\#})\right\}$ (23)
must hold (recall that $\Pi_{\mathcal{L}_{\#}^{\bot}}e_{i}(n)\neq 0$). To prove the opposite inequality, let $\mathcal{S}$ be an arbitrary element of $\mathbb{H}$, i.e., $\mathcal{S}\in\mathbb{J}(\mathcal{L}_{\#},\mathfrak{C}_{Het})$ and $T_{Het}$ is $\lambda_{\mu_{0}+\mathcal{S}}$-almost everywhere equal to a constant $C(\mathcal{S})$, say. Then Proposition B.1 in Appendix B of Pötscher and Preinerstorfer (2025) together with $\mathop{\mathrm{d}im}(\mathcal{L}_{\#})<n-1$ shows that $\mathcal{S}_{i}\subseteq\mathcal{S}$ holds for some $i\in I_{1}(\mathcal{L}_{\#})$. By Remark B.1(iv) given below, the condition in (10) is equivalent to
$e_{i}(n)\notin\mathsf{B}\text{ for every }i\in I_{1}(\mathcal{L}_{\#}).$
[…]
$C^{\ast}=\max\left\{T_{Het}(\mu_{0}+\Pi_{\mathcal{L}_{\#}^{\bot}}e_{i}(n)):i\in I_{1}(\mathcal{L}_{\#})\right\}.$
Using invariance of $T_{Het}$ w.r.t. addition of elements of $\mathcal{L}_{\#}$ (cf. Lemma A.2) we conclude that
$C^{\ast}=\max\left\{T_{Het}(\mu_{0}+e_{i}(n)):i\in I_{1}(\mathcal{L}_{\#})\right\}.$ (24)
Recall that $I_{1}(\mathcal{L}_{\#})\subseteq I_{1}(\mathfrak{M}_{0}^{lin})$. For $i\in I_{1}(\mathfrak{M}_{0}^{lin})\left\backslash I_{1}(\mathcal{L}_{\#})\right$. we have $i\in I_{0}(\mathcal{L}_{\#})$, and thus $e_{i}(n)\in\mathcal{L}_{\#}$. Since $\mathcal{L}_{\#}\subseteq\mathsf{B}$, $e_{i}(n)\in\mathsf{B}$ follows. Using Part 1 of Lemma A.2 and $\mathfrak{M}_{0}\subseteq\mathsf{B}$, we conclude that $\mu_{0}+e_{i}(n)\in\mathsf{B}$, and thus $T_{Het}(\mu_{0}+e_{i}(n))=0$. Since $T_{Het}$ is always nonnegative and since $I_{1}(\mathcal{L}_{\#})$ is nonempty, we can write (24) equivalently as
$C^{\ast}=\max\left\{T_{Het}(\mu_{0}+e_{i}(n)):i\in I_{1}(\mathfrak{M}_{0}^{lin})\right\}.$
The expression for $\alpha^{\ast}$ given in the theorem now follows immediately from the expression for $\alpha^{\ast}$ given in Part 1 of Theorem B.2.
Part 2-4 now follow from the corresponding parts of Theorem B.2 in light of what has been shown above. $\blacksquare$
Remark B.1: (Equivalent forms of the size-control conditions) (i) The proof of Lemma B.1 has shown that (15) is not only equivalent to (16), but also to (17) as well as to (18).
(ii) Non-inclusion statements of the form $\mathop{\mathrm{s}pan}\left(\left\{e_{i}(n):i\in J\right\}\right)\nsubseteqq\mathsf{B}$” ($J$ an index set) appearing in (15), (17), and (18) can equivalently be written as $e_{i}(n)\notin\mathsf{B}$ for some $i\in J$” due to the fact that $\mathsf{B}$ is a linear space (as $R$ is $1\times k$). Similarly, $\mathop{\mathrm{s}pan}\left(\left\{e_{i}(n):i\in J\right\}\right)\nsubseteqq\mathop{\mathrm{s}pan}(X)$” is equivalent to $e_{i}(n)\notin\mathop{\mathrm{s}pan}(X)$ for some $i\in J$”.
(iii) In the special case where $m=n$ and $n_{1}=n_{2}=...=n_{m}=1$, we learn from Lemma B.1 and its proof that (10) is equivalent to (19), as well as to (20). Since $\mathcal{I}_{\#}^{c}\subseteq I_{1}(\mathcal{L}_{\#})\subseteq I_{1}(\mathfrak{M}_{0}^{lin})$ by Part 3 of Lemma A.4, each one of (10), (19), and (20) is in turn equivalent to the condition
$e_{i}(n)\notin\mathop{\mathrm{s}pan}(X)\text{ for every }i\in I_{1}(\mathcal{L}_{\#}).$ (25)
[As a point of interest we note that conditions (10), (19), (20), and (25) are in fact equivalent also if, in the notation of Pötscher and Preinerstorfer (2025), we have $q\geq 1$, i.e., if a collection of $q$ restrictions is tested simultaneously. This can be seen by an inspection of the proofs of these equivalences. However, note that in case $q>1$ we have no result guaranteeing that these conditions are sufficient for size controllability of $T_{Het}$.]
(iv) Specializing Part (a) of Lemma B.1 and its proof to the case $n_{j}=1$ for $j=1,\ldots,n=m$, and noting that $\mathcal{I}_{\#}^{c}\subseteq I_{1}(\mathcal{L}_{\#})$ (Lemma A.4), one sees that further equivalent forms of (10) are given by the condition
$e_{i}(n)\notin\mathsf{B}\text{ for every }i\in I_{1}(\mathcal{L}_{\#}),$
as well as by the condition
$e_{i}(n)\notin\mathsf{B}\text{ for every }i\in\mathcal{I}_{\#}^{c},$
respectively. However, recall that while condition (7) implies anyone of the two equivalent conditions above, it is, in general, stronger in view of the examples in Appendix A.
(v) Since in the special case where $m=n$ and $n_{1}=n_{2}=...=n_{m}=1$ condition (10) appears also as the size-control condition for the standard (uncorrected) F-test statistic (see Pötscher and Preinerstorfer (2025)), this condition can also be written in any of the equivalent forms given in (iii) or (iv) in the case of testing a single restriction as considered here. [The equivalence of (10) with the other conditions in (iii) above even holds in the more general case where more than one restriction is subject to test.] We note that the before given equivalences do not rely on Assumption 1, an assumption that also does not appear in the size control results in Pötscher and Preinerstorfer (2025) for the classical (uncorrected) F-test statistic.
Remark B.2: The proof of Theorem 2.2 shows that $C^{\ast}$ defined in (9) can alternatively be written as in (24). The representation (24) has two advantages over (9): First, the index set $I_{1}(\mathcal{L}_{\#})$ is potentially smaller than $I_{1}(\mathfrak{M}_{0}^{lin})$ (see Lemma A.4); second, since $e_{i}(n)\notin\mathsf{B}$ for $i\in I_{1}(\mathcal{L}_{\#})$ under condition (10) (see Remark B.1(iv)), also $\mu_{0}+e_{i}(n)\notin\mathsf{B}$ for such $i$ ($\mu_{0}\in\mathfrak{M}_{0}$). Thus, (24) does not rely on the way $T_{Het}$ has been defined on the set $\mathsf{B}$.